Given { t, f, 7r, 3n, 9n,...}:

(a) Find the 12th term.

(b) Find the sum of the first 12 terms.



1
Expert's answer
2021-02-03T03:56:59-0500

(a)

"b_4=3n, b_5=9n,..."

"q=3"


"b_{12}=3^{12-3}n=19683n"

The 12th term is "19683n."


(b)

"b_4+b_5+...+b_{12}=3n\cdot\dfrac{1-3^9}{1-3}=29523n"

The sum of the first 12 terms is


"S_{12}=t+f+7r+29523n"